OpenStudy (anonymous):

write the product of 3x3x3x3x3 as a power

OpenStudy (vishweshshrimali5):

A power basically means multiplying the number that number of times. For example, \[\large{3^2 = 3 \times 3}\] \[\large{3^3 = 3\times3\times3}\] \[\large{3^4 = 3\times3\times3\times3}\]

OpenStudy (vishweshshrimali5):

Now can you tell the correct answer ?

OpenStudy (anonymous):

\[3^{5}\]

OpenStudy (anonymous):

yah @kritikachandak is right